# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v3

@@ 2.1. From Log-Sobolev to Transport
Let $\gamma$ be the standard Gaussian measure on $\mathbb{R}^{k}$. Let $X_{t}=e^{-t}X+\sqrt{1-e^{-2t}}Z$ be an Ornstein–Uhlenbeck process on $\mathbb{R}^{k}$ with distribution $\mu_{t}$, where $X\sim\mu$ and $Z\sim\gamma$ are independent. Then, the following hold.
  1. (1)
    For any $\mu\in P_{2}^{\mathrm{ac}}(\mathbb{R}^{k})$ satisfying (2.4) and any $t\geq 0$,
@@
  2. (2)
    For any $\mu\in P_{2}^{\mathrm{ac}}(\mathbb{R}^{k})$ and any $0\leq s\leq D(\mu\|\gamma)$, there is a probability measure $\pi$ such that $D(\pi\|\gamma)=s$ and
    $\sqrt{2s}+\mathsf{W}_{2}(\pi,\mu)\leq\sqrt{2D(\mu\|\gamma)}.$
For $s=0$, the inequality above reduces to the original version of Talagrand’s transport inequality.
@@
2.2. Generalization to $\Phi$-Sobolev
Theorem 6 can be generalized to the general $\Phi$-entropy setting (also known as $U$-functional or $f$-divergence). Let $\Phi:[0,+\infty)\rightarrow\mathbb{R}$ be a continuous convex function, twice differentiable on $(0,+\infty)$. The $\Phi$-entropy is defined as
$D_{\Phi}(\mu\|\nu):=\int\Phi(\rho)d\nu-\Phi(1).$
When $\rho$ is locally Lipschitz, the Fisher information associated with the $\Phi$-entropy, named the $\Phi$-Fisher information, [65] is defined as
$I_{\Phi}(\mu\|\nu):=\int\rho\Phi^{\prime\prime}(\rho)^{2}|\nabla\rho|^{2}{\text{\rm d}}\nu=\int\frac{|\nabla p(\rho)|^{2}}{\rho}{\text{\rm d}}\nu=\int\rho|\nabla\Phi^{\prime}(\rho)|^{2}{\text{\rm d}}\nu,$
where $p(r)=r\Phi^{\prime}(r)-\Phi(r)$. When $\Phi(t)=t\ln t$, the $\Phi$-entropy and the $\Phi$-Fisher information respectively reduce to the entropy and the Fisher information. The class of $\Phi$-entropy includes many important special entropies, which have found important applications in mathematics, information theory, statistics, and physics [65].
A common requirement for the function $\Phi$ is the following, which guarantees the displacement convexity of the $\Phi$-entropy [65, Chapter 16].
Definition 3 (Displacement convexity classes [54]).
The class $\mathcal{DC}_{\infty}$ is defined as the set of continuous convex functions $\Phi:\mathbb{R}_{+}\to\mathbb{R}$, twice continuously differentiable on $(0,+\infty)$, such that $\Phi(0)=0$, and, with the notation
$p(r)=r\Phi^{\prime}(r)-\Phi(r),\quad p_{2}(r)=rp^{\prime}(r)-p(r),$
$\Phi$ satisfies any one of the following equivalent differential conditions:
@@ 2.2. Generalization to $\Phi$-Sobolev
    $p_{2}\geq 0$;

### v4

@@ 2. From Log-Sobolev to Transport
Let $\gamma$ be the standard Gaussian measure on $\mathbb{R}^{k}$. Let $X_{t}=e^{-t}X+\sqrt{1-e^{-2t}}Z$ be an Ornstein–Uhlenbeck process on $\mathbb{R}^{k}$ with distribution $\mu_{t}$, where $X\sim\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathbb{R}^{k})$ and $Z\sim\gamma$ are independent. Then, for any $t\geq 0$,
@@
3. From Transport to Variant Isoperimetry
3.1. Problem Formulation and Preliminaries
We now consider a variant of the isoperimetric problem in which the perimeter of a set is identified by the size of an enlargement shell of the set. We then provide a solution to this variant isoperimetric problem by using the nonlinear transport inequality obtained in the last section.
Let $(\mathcal{X},d,\nu)$ be a Polish metric measure space with $\nu$ being a Borel probability measure. Denote $(\mathcal{X}^{n},d_{n},\nu_{n})$ as the $n$-fold product space of $\mathcal{X}$ equipped with the product metric $d_{n}$ and product measure $\nu_{n}=\nu^{\otimes n}$. Define the variant isoperimetric profile as for $a\in[0,1],t\geq 0$,
(3.1) $\Gamma_{n}(a,t):=\inf_{A:\nu^{\otimes n}(A)\geq a}\nu^{\otimes n}(A^{t}).$
We set $a=e^{-n\alpha}$ and $t=\sqrt{n\tau}$. Define the isoperimetric exponent as for $\alpha,\tau\geq 0$,
(3.2) $E_{n}(\alpha,\tau)$ $:=-\frac{1}{n}\ln\Gamma_{n}(e^{-n\alpha},\sqrt{n\tau}).$
Recall that the Wasserstein metric $\mathsf{W}_{2}(\pi_{X},\pi_{Y})$ (of order $2$) between two probability measures $\pi_{X}$ and $\pi_{Y}$ is defined by
(3.3) $\mathsf{W}_{2}^{2}(\pi_{X},\pi_{Y}):=\min_{\pi\in\Pi(\pi_{X},\pi_{Y})}\int d^{2}(x,y){\text{\rm d}}\pi(x,y),$
where $\Pi(\pi_{X},\pi_{Y})$ is the set of joint probability measures on $\mathcal{X}\times\mathcal{Y}$ with marginals equal to $\pi_{X},\pi_{Y}$ respectively. The existence of the minimizers are well-known; see, e.g., [72, Theorem 1.3]. For a probability measure $\pi_{W}$ and two transition probabilities $\pi_{X|W}$ and $\nu_{X|W}$, the conditional Wasserstein metric between $\pi_{X|W}$ and $\pi_{Y|W}$ given $\pi_{W}$ is defined by
$\mathsf{W}_{2}^{2}(\pi_{X|W},\pi_{Y|W}|\pi_{W}):=\int\mathsf{W}_{2}^{2}(\pi_{X|W=w},\pi_{Y|W=w}){\text{\rm d}}\pi_{W}(w).$
The conditional relative entropy of $\pi_{X|W}$ w.r.t. $\nu_{X|W}$ given $\pi_{W}$ is defined as
$D(\pi_{X|W}\|\nu_{X|W}|\pi_{W}):=\int D(\pi_{X|W=w}\|\nu_{X|W=w}){\text{\rm d}}\pi_{W}(w).$
Let $\nu_{X}$ and $\nu_{Y}$ be two probability measures on the space $(\mathcal{X},d)$. Define $E_{n}(\alpha,\tau|\nu_{X},\nu_{Y})$ similarly as the exponent in (3.2) but with $\Gamma_{n}$ replaced by
$\Gamma_{n}(a,t|\nu_{X},\nu_{Y}):=\inf_{A:\nu_{X}^{\otimes n}(A)\geq a}\nu_{Y}^{\otimes n}(A^{t}).$
Define
(3.4) $\psi_{\kappa}(\alpha,\tau|\nu_{X},\nu_{Y})$ $:=\sup_{\begin{subarray}{c}\pi_{XW}\in\mathscr{P}(\mathcal{X}\times\{0,1\}):\\ D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha\end{subarray}}\eta_{\tau,\kappa}(\pi_{XW}|\nu_{Y}),$
where
(3.5) $\eta_{\tau,\kappa}(\pi_{XW}|\nu_{Y})$ $:=\inf_{\begin{subarray}{c}\pi_{Y|XW}:\mathbb{E}[d^{2}(X,Y)]\leq\tau,\\ \mathbb{E}[\mathrm{Var}(d^{2}(X,Y)|X,W)]\leq\kappa\end{subarray}}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W}).$
Denote $\psi(\alpha,\tau|\nu_{X},\nu_{Y})=\psi_{\infty}(\alpha,\tau|\nu_{X},\nu_{Y})$, equivalently in which case, the variance constraint in (3.5) is removed. Without changing the optimal value, the range of $W$ in (3.4) can be changed to any larger finite set. This point can be proven by following proof steps similar to that of [77, Theorem 20].
Based on $\psi$ and $\psi_{\kappa}$, the asymptotic expression of $E_{n}$ for compactly supported $\nu_{X}$ is characterized in our previous work, which is a generalization of Ahlswede and Zhang’s work [3].
Theorem 7 (Asymptotic Exponent for Compactly-Supported Measures [77, Theorem 12]).
Let $(\mathcal{X},d)$ be a Polish metric space equipped with two probability measures $\nu_{X}$ and $\nu_{Y}$. The following hold.
@@ 3.1. Problem Formulation and Preliminaries
    Suppose $\nu_{X}$ is compactly supported. Then, for any $\alpha\geq 0,\tau\geq 0$,
    (3.6) $\limsup_{n\to\infty}E_{n}(\alpha,\tau|\nu_{X},\nu_{Y})\leq\lim_{\kappa\uparrow\infty}\psi_{\kappa}(\alpha,\tau|\nu_{X},\nu_{Y}).$
    In fact, a stronger result holds: Given $\delta,\kappa,\tau_{0}>0$, for sufficiently large $n$ and any set $A\subseteq\mathcal{X}^{n}$ such that $\nu_{X}^{\otimes n}(A)\geq e^{-n\alpha}$, there is a probability measure $\pi_{XW}\in\mathscr{P}(\mathcal{X}\times\mathbb{N})$ induced by $A$ such that $D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha+\delta$ and for any $\tau\in(\delta,\tau_{0}]$,
    $\frac{1}{n}\ln\nu_{Y}^{\otimes n}(A^{\sqrt{n\tau}})\leq(1+\delta)\eta_{\tau-\delta,\kappa}(\pi_{XW}|\nu_{Y})+\delta.$